OpenStudy (anonymous):

The perimeter of a regular hexagon is seventy-two and each side is represented by two x. What is x?

OpenStudy (anonymous):

The perimeter of a regular hexagon is 72. A regular hexagon means each side is the same. A hexagon is 6-sided figure. So if there is 6 sides and they all are the same length & add up to 72. Perimeter(72)/6(sides)= each side is 12. & each side is represented by 2x. We know each side's length is 12. So: 2x=12 /2 /2 x=6

OpenStudy (anonymous):

hope this helps ^^

OpenStudy (anonymous):

Would it be 12?

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

i see where you think 12 but when it gets to 12=2x, you have to get the x alone thus dividing 2 from both sides to get x=6 like this ^^: